Let the joint pmf of X and Y be defined by x+y 32 x 1,2, y,2,3,4 (a) Find fx(x), the marginal pmf of X. b) Find fyv), the marginal pmf of Y (c) Find P(XsY. (d) Find P(Y 2x). (e) Find P(X+ Y 3) (f) Find PX s3-Y) (g) Are Xand Y independent or dependent?Why or why not? (h) Find the means and the variances of X and Y
Answer #1

a) f(x) =\sum_{y=1}^{4}f(x,y) = (1/32)*((x+1)+(x+2)+(x+3)+(x+4))=(1/32)*(4x+10)=(2x+5)/16 for x=1,2

b)

f(y) =\sum_{x=1}^{2}f(x,y) =(1/32)*((1+y)+(2+y))=(2y+3)/32 for y=1,2,3,4

c)P(X>Y)=P(X=2,Y=1)=(2+1)/32=3/32

d)P(Y=2X)=P(X=1,Y=2)+P(X=2,Y=4)=(1+2)/32+(2+4)/32=9/32

e)P(X+Y=3)=P(X=1,Y=2)+P(X=2,Y=1)=(1+2)/32+(2+1)/32 =6/32=3/16

f)

P(X<=3-Y) =P(X=1,Y=1)+P(X=1,Y=2)+P(X=2,Y=1)=(1+1)/32+(1+2)/32+(2+1)/32=8/32=1/4

g)

as f(x,y) is not equal to f(x)*f(y) ' therefore X and Y are not independent

h)

from above marginal distribution of X:

x P(x) xP(x) x^2P(x)
1    7/16 0.4375 0.4375
2    9/16 1.1250 2.2500
total 1.0000 1.5625 2.6875
E(x) = 1.5625
E(x^2) = 2.6875
Var(x) E(x^2)-(E(x))^2 0.2461

hence E(X) =0.4375=25/16

Var(X)=0.2461 =63/256

similarly marginal pmf of Y:

y P(y) yP(y) y^2P(y)
1    5/32 0.1563 0.1563
2    7/32 0.4375 0.8750
3    9/32 0.8438 2.5313
4 11/32 1.3750 5.5000
total 1 2.8125 9.0625
E(y) = 2.8125
E(y^2) = 9.0625
Var(y) E(y^2)-(E(y))^2 1.1523

E(Y)=2.8125=45/16

Var(Y)=295/256